OpenStudy (anonymous):

This is the final part of a DE problem, i worked out the whole problem but i just do not know how i can solve for y in this case: (y/x)^2-(y/x)=ln(x)+C

OpenStudy (amistre64):

hmm, let v = y/x, or vx = y

OpenStudy (anonymous):

ok, i still dont see how i could solve it:\[v ^{2}-v=\ln(x)+C\]

OpenStudy (amistre64):

it might help to know how this all started

OpenStudy (anonymous):

Here is the original problem: Solve the DE:\[(x^2- xy + 2y^2 )dx + (x^2 - 2xy)dy = 0\]

OpenStudy (amistre64):

is it an exact DE?

OpenStudy (anonymous):

no, it fails the test

OpenStudy (amistre64):

bummer ... :) so your intent was to make it separable right?

OpenStudy (anonymous):

yeah, pretty much then just solve for y

OpenStudy (amistre64):

\[(x^2- xy + 2y^2 )dx + (x^2 - 2xy)dy = 0\] \[(x^2 - 2xy)dy = (x^2- xy + 2y^2 )dx\] \[(x - 2y)dy = (x- y + \frac{2}{x}y^2 )dx\] \[\frac{dy}{dx} = \frac{(x- y + \frac{2}{x}y^2 )}{(x - 2y)}\] \[y' = \frac{1- \frac yx + 2~(\frac yx)^2 )}{1 - 2\frac yx}\] \[y=vx~:~y'=v'x+v\] \[\frac{dv}{dx}x+v = \frac{1- v + 2~v^2}{1 - 2v}\] solve for dv/dx

OpenStudy (amistre64):

can we factor the right side any?

OpenStudy (amistre64):

\[\frac{dv}{dx}x+v = \frac{1- v + 2~v^2}{1 - 2v}\] \[\frac{dv}{dx}x = \frac{1- v + 2~v^2}{1 - 2v}-v\] \[\frac{dv}{dx}x = \frac{1- v + 2~v^2-v+2v^2}{1 - 2v}\] \[\frac{dv}{dx}x = \frac{1- 2v + 4~v^2}{1 - 2v}\] \[\frac{1 - 2v}{1- 2v + 4~v^2}{dv} = \frac 1xdx\] the left might be a good candidate for partial decomp ...

OpenStudy (psymon):

Isn't it possible to treat it like a quadratic? \[\frac{ y^{2} }{ x^{2} }-\frac{ y }{ x }=lnx + C\] \[y^{2}- xy - x^{2}(lnx + C) = 0\] Quadratic formula: \[\frac{ x \pm \sqrt{x^{2} - 4(-x^{2}lnx+Cx^{2}}) }{ 2 }\] ? Just an idea.

OpenStudy (anonymous):

oh, that makes sense, i think thats the correct approach

OpenStudy (psymon):

I just noticed it was quadratic in y, so seemed logical.

OpenStudy (amistre64):

im still trying to determine how it got to that form :)

OpenStudy (psymon):

Haha. Everything bu y is treated as a constant. So a is 1, b is -x, c is -x^2(lnx+C)

OpenStudy (amistre64):

but yes, the soluton from wolf is very quadratic formulaish looking

OpenStudy (psymon):

Might be able to just simplify it from where we have it then and accept that. Of course an initial condition would be nice so we don't have a C in there.

OpenStudy (anonymous):

yeah, its usually nice to get rid of the constant, but no initial condition was given in this problem.

OpenStudy (psymon):

Oh well, so we just deal with the C in there then and simplify it out : )

OpenStudy (psymon):

\[y = \frac{ x \pm \sqrt{x^{2} + 4x^{2}lnx +kx^{2}} }{ 2 }\] Where k = -4C